# HGT: Leveraging Heterogeneous Graph-enhanced Large Language Models for Few-shot Complex Table Understanding

### v1

@@ 4.3 Task-specific Instruction Tuning
Following the completion of the self-supervised task in Stage 1, the HGT has successfully aligned the representation spaces of the HG encoder and the LLM. This alignment, combined with the intrinsic capabilities of the LLM, empowers HGT to more effectively comprehend the topological nuances of complex tables. Consequently, when applied to a specific downstream task, HGT requires only a small number of training samples to quickly grasp the expected answer format for the current task and reorganize its pre-existing knowledge into the suitable output format. When fine-tuning in Stage 2, the model adopts the same weight-tuning strategy as in Stage 1, with all parameters frozen except for those of the HG Encoder.
@@ 5 Experiments
CTC TTC Table QA # Test Tables % Complex Tables # QA pairs Main Domains
@@
Model Type Models CTC (Macro–F1) TTC (Macro–F1) Table QA (Accuracy)
HiTab IM–TQA WCC IM–TQA WTQ (Flat) WTQ (Raw) HiTab IM–TQA
Bert-like ForTap 55.74 46.32 45.13 52.34 26.31 25.48 23.80 28.07
Encoder-based GetPt 57.45 48.61 47.53 55.42 22.63 22.33 22.29 21.05
Graph-based TabularNet 53.11 44.53 44.97 55.32 18.12 17.52 21.34 18.98
TabPrompt 64.44 45.75 49.24 53.32 16.62 13.72 19.19 13.08
Query Statement-based Binder 50.09 42.63 38.07 39.07
Aug-Codex 41.39 38.42 53.09 43.54
LLM-based TableLlama 58.58 52.67 46.31 54.69 37.85 33.32 47.05
GPT-3.5 62.46 55.63 52.32 59.10 39.25 37.51 42.05 51.52
HGT 66.24 60.11 56.10 59.58 45.00 43.11 54.61 52.95
Table 2: Overall evaluation results on three TU tasks with best bolded and runner-up underlined. ’–’ indicates that the current framework cannot handle the current task due to some limitations.  [^2] For a fair comparison with our framework, we use 3-shot prompting for Binder and Aug-Codex, which decreases performance compared to the results in their paper. More details are shown in Appendix C.
@@ 5.1 Datasets
To validate the effectiveness of our model, we selected a variety of datasets that are both widely studied and easy to parse, including TURL Deng et al. (2020), WCC Ghasemi-Gol and Szekely (2018), IM-TQA Zheng et al. (2023), HiTab Cheng et al. (2021b), WTQ (Flatten) Pasupat and Liang (2015), WTQ (Raw) Pasupat and Liang (2015), each pertinent to CTC, TTC, or Table QA, respectively.
Statistics for these datasets are presented in Table 1, illustrating variations in the types of annotations, the primary domains covered, and the proportion of complex tables. Given the focus on few-shot TU scenarios, we only list the number of tables within the test sets. Comprehensive details of the dataset annotations and pre-processing are available in Appendix A.
@@ 5.2 Baselines
We compare HGT with eight strong baselines to verify its effectiveness. The eight baselines can be categorized into four groups according to their frameworks. ForTapCheng et al. (2021a) and GetPtJia et al. (2023) emulate BERT, devising specific pretext tasks tailored for tabular data to pre-train the Transformer encoderDevlin et al. (2019). TabularNet Du et al. (2021) and TabPromptJin et al. (2023b) incorporate a graph encoder. BinderCheng et al. (2022) and Aug-CodexCao et al. (2023) rely on Codex to generate query statements to retrieve the answer. TableLlamaZhang et al. (2023) and GPT-3.5Ouyang et al. (2022b) are both LLMs renowned for their exceptional performance in few-shot scenarios.
@@ 5.3 Implementation Details
General Setup Across All Tasks. We employ Vicuna-7B-v1.5Chiang et al. (2023) as the base model for LLM and a 2-layer RGATWang et al. (2020a) as our tabular HG encoder whose hidden dimension is set as 1048. The initial vectors of nodes in HG are obtained by Sentence-BERT Reimers and Gurevych (2019), whose dimension is 768. We integrate a LoRAHu et al. (2021) module to the embedding layer of the LLM. At any tuning stage, HGT exclusively tunes the parameters of the RGAT and the LoRA module. This lightweight configuration allows our model to be efficiently trained on a single 4090 GPU. We assemble the tables from training and validation sets of the datasets mentioned above along with a selection of high-quality tables from the TURL as a comprehensive dataset consisting of 100k tables for use in the pre-training stage. Moreover, recognizing GPT’s limited sensitivity to numerical data, we prefix cells containing only numbers with "Float Value:" and so on, which can be achieved by regular expression. This enhancement is aimed at bolstering the semantic information conveyed by these cells.
General Setup across Evaluation Tasks. Given our aim to validate the HGT’s performance on few-shot TU, we randomly generate five 3-shot tasks for both training and validation on each dataset (i.e., 3 train/dev tables for CTC, 3 train/dev tables per class for TTC and 3 train/dev QA pairs for Table QA). Each training task, paired with a corresponding validation task, is utilized to fine-tune the models optimally for subsequent testing. Table 2 presents the average performance of a framework on the test set across these five tasks. In addition, to align the experiments more closely with real-world scenarios, we adhere to the setup outlined in Zheng et al. (2023) and refrain from explicitly providing the model with the hierarchical structure information in the header row / columns, despite some datasets being explicitly labeled with this detail. The detailed setup of baselines is shown in Appendix C.
Evaluation Metrics. For CTC and TTC, we adopt Macro-F1 as the evaluation metric. Considering that our framework, along with some baselines, employs a generative architecture, relying solely on exact match might inaccurately categorize some correct responses (e.g., generating an answer with an additional period at the end). Therefore, we incorporate the semantic-match Cheng et al. (2022) and human evaluation as the evaluation metric of Table QA.
@@ 5.4 Results and Analysis
The experimental results presented in Table 2 shows that HGT achieves the best performance on 7 out of 8 datasets, with the WTQ (Raw) dataset being the sole exception where Binder outperforms HGT. The following analysis sheds light on these outcomes: 1) While Query Statement-based frameworks demonstrate strong capabilities in Table QA, their architectural design limits their ability to perform other tabular tasks. Compared to other end-to-end frameworks, their generality is somewhat lacking. 2) WTQ (Raw) represents the unmodified version of WTQ (Flat), retaining the complexity of hierarchical headers within tables. A visual depiction of the difference between these versions is provided in Fig. 8. The noticeable drop in performance of other methods on WTQ (Raw) underscores HGT’s capability to effectively process complex table structures. 3) The fact that other LLM-based frameworks underperform compared to HGT underscores the efficacy of our framework in mitigating the drawbacks associated with linearized table representations which locks the full capabilities of LLMs. 4) Without sufficient data for fine-tuning the downstream tasks, pre-trained models fail to reach their full potential. 5) On the IM-TQA dataset, the performance of Query Statement-based frameworks falls short compared to the LLM-based frameworks. We surmise this discrepancy is likely due to the dataset featuring particularly complex tables and look-up type questions, while Statement-based frameworks excel in different problem types, such as arithmetic queries. The results of the experiments exploring HGT at different numbers of training samples are detailed in Appendix. D.
@@ 5.5 Ablation Study
We conduct ablation experiments on a single dataset IM-TQA, the only dataset containing three types of labels, to validate the effectiveness of each component and self-supervised objective, the results of which are shown in Table 3.
We first conduct ablation studies $w/o$ three self-training tasks respectively. We pre-train three models without TRC, TCM, and TCG respectively to assess their individual impact. These three pre-training tasks prove beneficial across all table evaluation tasks, with TCM offering the most substantial enhancement across three evaluation tasks—enhancements of 5.54%, 3.53%, and 6.70%. This improvement is akin to how a human can only effectively organize disordered cells by understanding the table’s semantics and the interrelations between cells. The contribution of TRC is particularly noticeable in CTC and Table QA compared to TTC, with improvements of 4.53% and 5.95% respectively. as it aids the model in recognizing that the semantics of cells in header rows and data rows differ, even when their content is identical. TCG is especially advantageous for TTC, more so than the other tasks, aligning with our expectations.
We then conduct ablation studies about our method to construct tabular graphs. This involves converting tables into homogeneous graphs rather than heterogeneous ones and replacing the heuristic linking strategy with a raw method linking all neighboring cells. Both variants result in substantial performance declines in CTC and Table QA, with at least 3.15% and 3.51% reduction on these two tasks, respectively. It indicates that our conversion method is more effective than existing methods for modeling complex tables.
Modules CTC TTC Table QA
Full HGT 66.24 59.58 52.95
$\hskip 4.26773ptw/o$ TRC 61.71 58.03 47.05
$\hskip 4.26773ptw/o$ TCM 60.70 56.05 46.25
$\hskip 4.26773ptw/o$ TCG 65.38 57.46 50.56
$\hskip 4.26773ptw/o$ HG 62.42 58.55 49.44
$\hskip 4.26773ptw/o\ hl$ 63.09 58.87 48.96
Table 3: Ablation results on IM-TQA dataset. The "$w/o$ TRC", "$w/o$ TCM", and "$w/o$ TCG": HGT pre-trained without the specified objective. The "$w/o$ HG": HGT pre-trained with homogeneous graph. The "$w/o\ hl$": HGT pre-trained without the heuristic linking strategy described earlier, instead linking all neighboring cells directly.
@@ 6 Conclusion
In this paper, we introduce a novel framework, HGT, tailored for few-shot complex TU. The effectiveness of HGT is validated across multiple datasets for CTC, TTC, and Table QA, accompanied by an in-depth ablation study to examine the impact of each component.
In future work, we plan to expand HGT’s applicability to tables featuring more diverse layouts and to further improve HGT’s performance in Table QA tasks by implementing techniques that enhance the model’s inference capabilities.
Limitations
Firstly, HGT’s performance notably declines when dealing with tables that have irregular layouts, such as those containing sub-titles or images, or forms filled with personalized information. Secondly, HGT tends to consume excessive GPU memory when processing larger tables. Lastly, when faced with Excel-type forms, it is needed to use the API to convert them to html-format tables first.
Appendix A Datasets and Pre-processing
Figure 5: Annotation correspondence of HiTab.
Figure 6: Manual Features for Table Orientation Classification.
IM-TQA Zheng et al. (2023) consists of tables in Chinese, gathered from encyclopedias and reports, and translated into English. These tables are annotated for the three tasks: CTC, TTC, and Table QA. Due to our method of converting tables into HGs not being suitable for Vertical tables—where header cells do not align with the table rows—we employ a feature engineering method to pre-process and convert Vertical tables into Horizontal tables. This method has demonstrated a 98 % accuracy rate, with the relevant features depicted in Fig. 6. HiTab Cheng et al. (2021b) features hierarchical matrix tables sourced from statistical reports and Wikipedia articles, annotated for row and column granularity as well as Table QA. We also transform these broad-granularity annotations into cell-level annotations, as shown in Fig. 5. WTQ (Flatten) and WTQ (Raw) originate from the same work Pasupat and Liang (2015) and consist of tables sourced from Wikipedia, accompanied by questions that involve basic arithmetic operations like sum and max, and sometimes necessitate compositional reasoning. To simplify the Table QA task, the authors converted the hierarchical structure of header rows into a single row in the WTQ (Raw) dataset, resulting in the WTQ (Flatten) version. An illustration of this transformation is presented in Fig. 8. WCC Ghasemi-Gol and Szekely (2018), a subset of the July 2015 Common Crawl, is annotated for TTC using the web table taxonomy introduced by Crestan and Pantel 2011. It is not explicitly divided into training and test sets, so we use the 353 tables that were not sampled as the test set. Due to the GPU memory limit, we exclude the tables containing more than 150 cells.
Appendix B Top Header Row Number
The heuristic rule of acquiring the Top Header Row Number is shown in Alg. 1. Alg. 2 aims to determine the type of each row of a table.
Algorithm 1 Acquisition of Top Header Row Number
1:function $getThrn$(Table $X$)
2:     TopHeadRowNumber $thrn_{X}\leftarrow 0$
3:     for Cell $x$ in $X.firstRow$ do:
4:         // row span is an attribute of each cell
5:         RowSpan $rs\leftarrow$ x.getRowSpan()
6:         $thrn\leftarrow max(rs,thrn)$
7:     end for
8:     Return $thrn$
9:end function
Algorithm 2 Identify Row Type
1:function $identifyRowType$(Table $X$, Row $R$)
2:     $thrn_{X}\leftarrow getThrn(X)$
3:     $r_{th}\leftarrow getRowLocationNumber(X)$
4:     // $r_{th}$ represents $R$ located in the $r_{th}$ row of table $X$
5:     Cell $x\leftarrow R.getFirstCell()$
6:     // Cell $x$ is the first cell located in row $R$
7:     if $r_{th}<thrn_{X}$ then
8:         Return Header-Row
9:     else if $x.getColumnSpan()=X.getColumnNumber$ then
10:         Return Header-Row
11:     else
12:         Return Data-Row
13:     end if
14:end function
Appendix C Setup of Baselines
This section introduces the setup of baselines. GetPt and Aug-Codex have not made their source code publicly available as of yet, leading us to replicate the framework based on the details provided in their paper. TableLlama, which does not utilize the Cell Type Classification (CTC) and Table Type Classification (TTC) datasets for training, exhibits suboptimal performance when inputs are presented alongside in-context prompts during testing. To address this, we fine-tuned the model’s parameters using few-shot training data prior to testing. Additionally, given that Tablellama was trained using a substantial volume of data from the HiTab dataset, its test results are not indicative of performance in few-shot scenarios. For GPT-3.5, we employed API calls for the testing, where the training data was integrated as in-context prompts during the test phase.
Appendix D Results of varied training shots on HiTab
We further explore the performance of HGT under different numbers of training samples with the results depicted in Fig. 7. We choose Aug-codex, Fortap, and Tablellama as the baselines for comparison. In Fig. 7, the vertical axis represents the accuracy, and the horizontal axis represents the number of training samples, which are 1, 3, 5, and full, respectively. The full setting refers to the use of any amount of data from the training set until optimal results are achieved by the model. It is also noteworthy that Tablellama’s accuracy is only reported in full settings, as it incorporates HiTab training data during its training phase.
As you can see from the graph, HGT performs best with a limited number of samples and maintains competitiveness with the state-of-the-art Aug-Codex even in full settings.
Figure 7: Impact of varied shots on HiTab for Table QA.
Figure 8: Comparison of a table before and after being flattened
Figure 9: Examples of three downstream table-related tasks presented as instructions include: Cell Type Classification, Table Type Classification, and Table QA.
Appendix E Downstream Table-related Task Examples
Sample inputs for CTC, TTC, and Table QA are shown in Fig. 9.

### v2

@@ 4.3 Stage 2: Task-specific Instruction Fine-tuning
Following the completion of the self-supervised pre-training tasks in Stage 1, the HeGTa has successfully aligned the representation spaces of the HG encoder and the LLM. This alignment, combined with the intrinsic capabilities of the LLM, empowers HeGTa to effectively comprehend the topological nuances of complex tables. Consequently, when applied to a specific downstream task (i.e., CTC, TTC, TQA), HeGTa requires only a small number of training samples to grasp the expected answer format for the downstream task and reorganize its pre-existing knowledge into the suitable output format. Example prompts of downstream tasks can be found in Appendix. When fine-tuning in Stage 2, the model adopts the same tuning strategy as in Stage 1, with all parameters frozen except for those of the HG Encoder. To enhance the reasoning capabilities of HeGTa, we annotated the reasoning process with Chain of Thought (CoT) (Wei et al. 2022) for $k$-shot TQA samples. Inspired by Auto-CoT (Zhang et al. 2022), we first clustered the table samples in the test set, and then clustered the corresponding questions within each table cluster. Finally, we selected the central TQA samples from each question cluster for annotation. This approach ensures the diversity of contextual examples with a limited number of annotations.
@@ 5 Experiments
CTC TTC TQA # Test Tables % Complex Tables # QA pairs Main Domains
@@
Model Type Models CTC TTC TQA
HiTab IM–TQA WCC IM–TQA WTQ (Flat) WTQ (Raw) TabFact HiTab IM–TQA
Bert-like ForTap 56.33 47.34 46.14 53.40 27.33 26.52 28.67 24.88 29.08
Encoder-based GetPt 58.45 49.67 48.62 56.45 23.69 23.34 25.03 23.34 22.07
Graph-based TabularNet 54.14 45.62 46.01 56.41 19.18 18.58 20.81 22.34 20.05
TabPrompt 63.56 46.77 50.31 54.34 17.71 14.77 19.56 20.20 14.14
LLM-based (Symbolic) Binder 43.16 40.05 59.04 44.11 41.25
Aug-Codex 33.32 32.28 57.07 45.87 42.16
COTable 44.38 42.97 59.76 49.68 43.06
LLM-based (Textual) TableCoT 53.12 49.76 45.57 53.80 31.42 32.12 52.01 47.55 45.21
TableLlama 60.15 53.40 51.56 57.55 38.88 39.50 48.82
HeGTa 66.39 59.58 55.32 58.46 45.12 45.96 60.61 51.56 50.73
Table 2: Overall evaluation results on three TU tasks with best bolded and runner-up underlined. ‘–’ indicates that the current framework cannot handle the current task due to some limitations.  [^2] For a fair comparison, all LLM-based methods use the open-source LLaMA2-7B (Touvron et al. 2023) as the underlying model. More details of baselines are shown in Appendix.
@@ 5.1 Datasets
To evaluate HeGTa, we selected a variety of datasets that are widely studied and easy to parse, including WCC (Ghasemi-Gol and Szekely 2018), IM-TQA (Zheng et al. 2023), HiTab (Cheng et al. 2021b), WTQ (Raw) along with its header-flatten version WTQ (Flatten) (Pasupat and Liang 2015) and TabFact (Chen et al. 2020), each pertinent to CTC, TTC, or TQA, respectively.
Statistics for these datasets are presented in Table 1, illustrating the types of annotations, the primary domains covered, and the proportion of complex tables. Given the focus on few-shot TU scenarios, we only list the size of test sets. Comprehensive details of the dataset annotations and pre-processing are available in Appendix.
@@ 5.2 Baselines
We compare HeGTa with several strong baselines to verify its effectiveness. These baselines can be categorized into four groups according to their frameworks. ForTap (Cheng et al. 2021a) and GetPt (Jia et al. 2023) emulate BERT, devising specific pretext tasks tailored for tabular data to pre-train the Transformer encoder (Devlin et al. 2019). TabularNet (Du et al. 2021) and TabPrompt (Jin et al. 2023b) incorporate a graph encoder. Binder (Cheng et al. 2022), Aug-Codex (Cao et al. 2023) and COTable (Wang et al. 2024) leverage LLMs to generate symbolic programs, which are then executed to obtain the final answer. Compared to symbolic methods, TableLlama (Zhang et al. 2023a) and TableCoT (Chen 2022) leverage the textual understanding capabilities of LLMs to directly address tabular tasks.
@@ 5.3 Implementation Details
General Setup Across All Tasks. We employ Vicuna-7B-v1.5 (Chiang et al. 2023) as the base LLM and a 2-layer Heterogeneous Graph Transformer (HGT) (Wang et al. 2020a) as our tabular HG encoder whose hidden dimension is set as 1024. The initial vectors of nodes in HG are obtained by S-BERT (Reimers and Gurevych 2019), whose dimension is 768. We integrate a LoRA (Hu et al. 2021) module to the embedding layer of the LLM. At any tuning stage, HeGTa exclusively tunes the parameters of the HGT and the LoRA module. This lightweight configuration allows our model to be trained on a single 4090 GPU. We assemble the tables from training and validation sets of the datasets mentioned above along with a selection of TURL (Deng et al. 2020) corpus as a comprehensive dataset consisting of 200k tables for use in the pre-training stage.
Module CTC TTC TQA
HiTab IM–TQA WCC IM–TQA WTQ (Flat) WTQ (Raw) TabFact HiTab IM–TQA
$\hskip 4.26773ptw/o$ TRC -2.65 -1.53 +1.02 -1.40 -1.10 -1.34 -0.89 -0.78 +1.19
$\hskip 4.26773ptw/o$ TCM -4.50 -4.88 -3.22 -2.44 -2.31 -2.71 -2.10 -1.67 -2.45
$\hskip 4.26773ptw/o$ TCG +0.11 +1.19 -1.61 -1.73 -2.21 -2.54 -2.57 -0.30 -2.58
$\hskip 4.26773ptw/o$ HG -3.28 -1.61 -0.24 -0.60 -1.80 -1.75 -1.68 -1.01 -1.31
$\hskip 4.26773ptw/o$ hl -2.25 +0.90 +0.42 -0.75 -0.74 -0.85 -0.79 -0.20 +0.56
$\hskip 4.26773ptw/o$ CoT -2.31 -2.35 -2.57 -1.98 -1.31
Table 3: Ablation results on all datasets. “$w/o$ TRC”, “$w/o$ TCM”, and “$w/o$ TCG”: HeGTa pre-trained without the specified objective. “$w/o$ HG”: HeGTa pre-trained with homogeneous graphs. “$w/o\ hl$”: HeGTa pre-trained with linking all neighboring cells directly instead of the heuristic linking strategy.
General Setup across Evaluation Tasks. Following a typical $k$-shot setup (Liu et al. 2021; Wang et al. 2020b) to validate the HeGTa’s performance on few-shot TU, we randomly generate five 8-shot tasks for both training and validation on each dataset (i.e., 8 train/dev tables for CTC, 8 train/dev tables per class for TTC and 8 train/dev QA pairs for TQA). Each training task, paired with a corresponding validation task, is utilized to fine-tune the models optimally for subsequent testing. Table 2 presents the average performance on the test set across the five tasks.
Evaluation Metrics. We adopt Macro-F1, Macro-F1 and accuracy as the evaluation metric for CTC, TTC and TQA, respectively. Given the generative structure of LLMs, relying solely on exact matches might inaccurately categorize some correct responses (e.g., adding an answer with an additional period at the end). Therefore, we incorporate the semantic-match evaluator (Cheng et al. 2022) for TQA, which preprocesses the outputs prior to evaluation.
@@ 5.4 Results and Analysis
Table 2 compares the performance of different methods on all datasets described in § 5.1. From the results, We have the following observations: 1) HeGTa outperformed all baselines across 9 datasets, demonstrating its superiority over both non-LLM methods and LLMs with equivalent parameters. Additionally, as shown in Appendix, HeGTa demonstrates competitive performance in embedding-related tasks (i.e., CTC and TTC) even compared to SOTA GPT-4. 2) LLM-based symbolic methods excel in TQA but are inherently incapable of handling other table tasks due to their architectural design. Moreover, methods with 7B parameters struggle with reasoning and generating precise, executable code compared to those based on ChatGPT. For further details, see Appendix. 3) WTQ (Raw) represents the unprocessed version of WTQ (Flat), preserving the complexity of hierarchical table headers. The significant performance drop of other methods on WTQ (Raw) highlights HeGTa’s ability to effectively handle relational tables with complex structures. 4) Compared to HeGTa, other LLM-based frameworks underperform due to the loss of table topological information, even with enhancements like TableLlama’s extensive training on linearized tables or TableCoT’s use of HTML, a format better suited for LLMs (Sui et al. 2024). This underscores the effectiveness of introducing GNNs in overcoming the limitations of linearized table representations, thereby unlocking the full potential of LLMs.
Strengths and Weaknesses of LLM-Based Models
All QA samples of HiTab can be classified into lookup-based reasoning (L-R) and computation-based reasoning (C-R) (e.g., sum, max, count). Table 4 presents the performance of three LLM-based methods on two types of QA samples. TableCoT shows superior performance in L-R compared to C-R, aligning with the conclusion (Chen et al. 2022) that LLMs struggle with numerical calculations. COTable leverages SQL to mitigate the shortcomings of LLMs in numerical reasoning, leading to the highest accuracy in C-R. However, the hierarchical structure causes a decline in SQL performance. HeGTa addresses the issue of structural information loss of linearized tables. As a result, HeGTa achieves notable improvements, particularly in L-R samples. Additionally, without restricting to the few-shot setting, i.e., tuning the model with the entire training set, HeGTa consistently maintains top performance, as shown in Appendix.
L-R C-R Total
# Test set 1133 100.00 451 100.00 1584 100.00
TableCoT 612 54.02 195 43.24 806 47.55
COTable 578 51.02 265 58.76 842 49.68
HeGTa 669 59.05 206 45.68 874 51.56
Table 4: Further results of the three methods on HiTab. “#” represents the number of correctly answered samples, and “%” denotes the proportion of correctly answered samples.
@@ 5.5 Ablation Study
We conduct ablation experiments to evaluate each component’s effectiveness, as shown in Table 3.
We begin by examining the individual improvements contributed by each of the three self-training tasks. To evaluate their specific impact, we pre-train three models, each omitting one of the objectives: TRC, TCM, or TCG. These three pre-training objectives improve the LLM’s understanding of tabular structure. The TCM task most effectively aligns the tabular modality with the LLM’s semantic space. The TRC task enhances performance on relational tables, while the TCG task boosts TQA performance but slightly reduces CTC performance.
We conducted ablation studies on CoT and our tabular graph construction method by removing the CoT text, converting tables into homogeneous rather than heterogeneous graphs, and replacing the heuristic linking (hl) strategy with a naive rule that links all neighboring cells. The results clearly demonstrate the significant impact of CoT on TQA tasks. The introduction of HGs benefits all tasks, with particularly notable improvements in TU with hierarchical structures. However, compared to the benefits of HGs, the hl strategy shows less pronounced gains for non-relational tables, possibly because the naive rule is more effective for aggregating and transmitting information in these tables.
@@ 6 Conclusion
In this paper, we introduce a novel framework, HeGTa, tailored for few-shot complex TU. The effectiveness of HeGTa is validated across multiple tabular datasets, accompanied by an in-depth ablation study to examine the impact of each component. In future work, we plan to expand HeGTa’s applicability to tables featuring more diverse layouts and to further improve HeGTa’s performance in TQA tasks by integrating the program-aided paradigm into our approach.
7 Acknowledgments
This work is partially supported by National Nature Science Foundation of China under No. 62476058, by the project ”Key Laboratory of rich-media Digital Publishing Content Organization and Knowledge Service Open Fund-Research on Knowledge-enhanced Training Techniques of Large Language Model” No. ZD2024-04/01. We thank the Big Data Computing Center of Southeast University for providing the facility support on the numerical calculations in this paper.
[…]